Hi I have a few problems that I am just stuck on! I appreciate any assistance that you can offer...I am terrible at alegbra and have no clue to help understand it. time is of essence...thanking you in advance for your time and understanding.
Here are the problems:
Find the slope and the y- intercept of the line. Write the fractional answer in lowest terms.
y= -7/4x - 1/9 the slope is:??? the y intercept is(0, )
2. find the slope and the y intercept:
y=2.9x - 5 the slope is:??? the y intercept is (0,)
3. 3x = 4y + 24 (ditto as above)
4. 7- 1/5 y = 2x
I hope you can help...

Answer by Alan3354(69443)

Find the slope and the y- intercept of the line. Write the fractional answer in lowest terms.
y= -7/4x - 1/9 the slope is:??? the y intercept is(0, )
This is in slope-intercept form, y = mx + b
m = slope
b is the y-intercept
m = -7/4 slope
b = -1/9 --> (0,-1/9)
----------------------------
2. find the slope and the y intercept:
y=2.9x - 5 the slope is:??? the y intercept is (0,)
Same as above
--------------------
3. 3x = 4y + 24 (ditto as above)
Solve for y.
--------------
4. 7- 1/5 y = 2x
Solve for y.